# On $f$-Divergence Principled Domain Adaptation: An Improved Framework

### v1

@@ C.1 $\chi^{2}$-Divergence
If we further consider the affine transformation of Lemma 2.1 and the scaling transformation of Lemma 2.2, we can recover Hammersley-Chapman-Robbins lower bound and the Cramér-Rao and van Trees lower bounds (Polyanskiy & Wu, 2023).
@@ C.3 Jeffereys Divergence
Moreover, there is a tight relationship between Jensen-Shannon (JS) divergence and Jeffreys divergence (Crooks, 2008): $\mathrm{D}_{\rm JS}(P||Q)\leq\left\{\frac{1}{4}\mathrm{D}_{\rm Jeffereys}(P||Q),\log{\frac{2}{1+e^{-\mathrm{D}_{\rm Jeffereys}(P||Q)}}}\right\}$.Although we don’t directly use JS divergence in our algorithms, minimizing Jeffreys divergence simultaneously minimizes JS divergence.
@@ Appendix D Omitted Proofs and Additional Results
Let $f_{\nu}$ and $f_{\mu}$ be the ground truth labeling functions for the target domain and source domain, respectively, i.e. $f_{\nu}(x)=\arg\max_{y\in\mathcal{Y}}\nu(Y=y|x)$ and $f_{\mu}(x)=\arg\max_{y\in\mathcal{Y}}\mu(Y=y|x)$.
@@
where Eq. (11) and (13) are by the triangle property, Eq. (12) is by Lemma B.2, and the last inequality is by the definition of $f$-DD. This completes the proof. ∎
@@
$R_{\nu}(h)\leq$ $*\displaystyle R_{\mu}(h)+\widetilde{\mathrm{D}}^{h,\mathcal{H}}_{\phi}(\mu||\nu)+\lambda^{*}$
$\leq$ $*\displaystyle R_{\hat{\mu}}(h)+2\hat{\mathfrak{R}}_{\mathcal{S}}(\mathcal{H}^{\ell})+\mathcal{O}\left(\sqrt{\frac{\log(1/\delta)}{n}}\right)+\widetilde{\mathrm{D}}^{h,\mathcal{H}}_{\phi}(\mu||\nu)+{\lambda}^{*}$ (18)
@@
where Eq. (18) and Eq. (19) are by Lemma B.3 and Lemma D.1, respectively. For the last inequality, notice that the loss is bounded between $[0,1]$ so the exponential function is $e$-Lipschitz in this domain, then we apply the Talagrand’s lemma (Mohri et al., 2018, Lemma 5.7), i.e. ^$\hat{\mathfrak{R}}_{\mathcal{T}}(\exp\circ\mathcal{H}^{\ell})\leq e\hat{\mathfrak{R}}_{\mathcal{T}}(\mathcal{H}^{\ell})$.
Finally, due to the boundedness of the loss, there always exists such a constant $\beta\in(0,e]$, and a simple choice is $\beta=1$. This completes the proof. ∎
D.3 Proof of Lemma 4.1
@@
We hint that if the above step is handled more carefully, then one can prove a bound for the absolute mean deviation (i.e. $\left|A\right|$) in the end. The current development is enough for our purpose.
@@
$*.R_{\nu}(h)-R_{\mu}(h)\leq\inf_{t\geq 0}\frac{1}{t}\left(K_{\mu}(t)+\mathrm{D}^{h,\mathcal{H}}_{\phi}(\nu||\mu)\right)+\lambda^{*}.$
@@
$R_{\nu}(h)-R_{\mu}(h)\leq$ $*.\displaystyle\inf_{t}\frac{t}{2\phi^{\prime\prime}(1)}+\frac{\mathrm{D}^{h,\mathcal{H}}_{\phi}(\nu||\mu)}{t}+\lambda^{*}=\sqrt{\frac{2\mathrm{D}^{h,\mathcal{H}}_{\phi}(\nu||\mu)}{\phi^{\prime\prime}(1)}}+\lambda^{*}.$
@@
D.6 Proof of Lemma 4.2 and Generalization Bounds for $f$-DD
@@
$=$ $\sup_{h^{\prime},\alpha}{{{\mathbb{E}_{\nu}}\left[t_{0}\ell(h,h^{\prime})\right]}+{{\mathbb{E}_{\mu}}\left[\alpha-\phi^{*}(t_{0}\ell(h,h^{\prime})+\alpha)\right]}}-\sup_{h^{\prime},t,\alpha}{{\mathbb{E}_{\hat{\nu}}}\left[t\ell(h,h^{\prime})\right]}+{{\mathbb{E}_{\hat{\mu}}}\left[\alpha-\phi^{*}(t\ell(h,h^{\prime})+\alpha)\right]}$
$\leq$ $\sup_{h^{\prime},\alpha}{{{\mathbb{E}_{\nu}}\left[t_{0}\ell(h,h^{\prime})\right]}+{{\mathbb{E}_{\mu}}\left[\alpha-\phi^{*}(t_{0}\ell(h,h^{\prime})+\alpha)\right]}-\left({{\mathbb{E}_{\hat{\nu}}}\left[t_{0}\ell(h,h^{\prime})\right]}+{{\mathbb{E}_{\hat{\mu}}}\left[\alpha-\phi^{*}(t_{0}\ell(h,h^{\prime})+\alpha)\right]}\right)}$
$\leq$ $\sup_{h^{\prime}}\left|{{\mathbb{E}_{\nu}}\left[t_{0}\ell(h,h^{\prime})\right]}-{{\mathbb{E}_{\hat{\nu}}}\left[t_{0}\ell(h,h^{\prime})\right]}\right|+\sup_{h^{\prime},\alpha}\left|{{\mathbb{E}_{\mu}}\left[\phi^{*}(t_{0}\ell(h,h^{\prime})+\alpha)\right]}-{{\mathbb{E}_{\hat{\mu}}}\left[\phi^{*}(t_{0}\ell(h,h^{\prime})+\alpha)\right]}\right|$
@@
where the last two inequalities are by the scaling property of Rademacher complexity and Talagrand’s lemma (Mohri et al., 2018). This completes the proof. ∎
@@
Under the conditions in Lemma 4.2. Let $\phi$ be twice differentiable and ϕ′′superscriptitalic-$\phi^{\prime\prime}$ is monotone, for any $h\in\mathcal{H}$, with probability at least $1-\delta$, we have
@@
Then, recall that $\ell\leq 1$, and we use Lemma B.5 to obtain that
@@
$=$ $B(t)t^{2}{{\mathbb{E}_{\mu}}\left[(g(X))^{2}\right]}+B(t)t^{2}(C^{\prime 2}-2C^{\prime})\left({{\mathbb{E}_{\mu}}\left[g(X)\right]}\right)^{2}-tC{{\mathbb{E}_{\mu}}\left[g(X)\right]}$
$\leq$ $B(t)t^{2}{{\mathbb{E}_{\mu}}\left[(g(X))^{2}\right]}+B(t)t^{2}C^{2}\left({{\mathbb{E}_{\mu}}\left[g(X)\right]}\right)^{2}-tC{{\mathbb{E}_{\mu}}\left[g(X)\right]},$ (24)
@@
Since ${{\mathbb{E}_{\mu}}\left[g(X)\right]}\leq\min\{r_{1}+r,1\}$ and $g(\cdot)\leq 1$, we have
$B(t)t{{\mathbb{E}_{\mu}}\left[(g(X))^{2}\right]}+B(t)tC^{2}\left({{\mathbb{E}_{\mu}}\left[g(X)\right]}\right)^{2}-C{{\mathbb{E}_{\mu}}\left[g(X)\right]}$
$\leq$ $B(t)t{{\mathbb{E}_{\mu}}\left[(g(X))\right]}+B(t)tC^{2}\min\{r_{1}+r,1\}{{\mathbb{E}_{\mu}}\left[g(X)\right]}-C{{\mathbb{E}_{\mu}}\left[g(X)\right]}.$
@@
$B(t)t{{\mathbb{E}_{\mu}}\left[g(X)\right]}+B(t)tC^{2}\min\{r_{1}+r,1\}{{\mathbb{E}_{\mu}}\left[g(X)\right]}-C{{\mathbb{E}_{\mu}}\left[g(X)\right]}\leq 0.$
@@
$\left(e^{t}-t-1\right)\left(1+C^{2}\min\{r_{1}+r,1\}\right)-tC\leq 0,$ (25)
@@
$R_{\nu}(h)\leq R_{\hat{\mu}}$ $(h)+\inf_{C_{1},C_{2}}\frac{\mathrm{D}^{h,\mathcal{H}_{r}}_{\rm KL}(\hat{\nu}||\hat{\mu})}{C_{1}}+C_{2}R^{r}_{\mu}(h)+\mathcal{O}\left(\hat{\mathfrak{R}}_{\mathcal{S}}\left(\mathcal{H}^{\ell}_{r_{1}}\right)+\hat{\mathfrak{R}}_{\mathcal{S}}\left(\mathcal{H}^{\ell}_{r}\right)\!+\!\hat{\mathfrak{R}}_{\mathcal{T}}\left(\mathcal{H}^{\ell}_{r}\right)\right)+\mathcal{O}\left(\frac{\log(1/\delta)}{n}\!+\!\frac{\log(1/\delta)}{m}\right)$
$+\mathcal{O}\left(\sqrt{\frac{r_{1}\log(1/\delta)}{n}}+\sqrt{\frac{r\log(1/\delta)}{n}}+\sqrt{\frac{r\log(1/\delta)}{m}}\right)+\lambda_{r}^{*}.$
@@
D.10 Generalization Bounds based on $\chi^{2}$-DD

### v2

@@ C.1 $\chi^{2}$-Divergence
If we further consider the affine transformation of Lemma 2.1 and the scaling transformation of Lemma 2.2, we can recover Hammersley-Chapman-Robbins lower bound and the Cramér-Rao and van Trees lower bounds [30].
@@ C.3 Jeffereys Divergence
Moreover, there is a tight relationship between Jensen-Shannon (JS) divergence and Jeffreys divergence [73]: $\mathrm{D}_{\rm JS}(P||Q)\leq\left\{\frac{1}{4}\mathrm{D}_{\rm Jeffereys}(P||Q),\log{\frac{2}{1+e^{-\mathrm{D}_{\rm Jeffereys}(P||Q)}}}\right\}$.Although we don’t directly use JS divergence in our algorithms, minimizing Jeffreys divergence simultaneously minimizes JS divergence.
C.4 Understanding the Importance of a Tighter Variational Representation
In simplified terms, Lemma 2.1 gives a variational representation expressed as:
$\mathrm{D}_{\phi}(P||Q)=\sup_{g\in\mathcal{G}}V_{1}(g),$
while Lemma 2.2 provides a variational representation as:
$\mathrm{D}_{\phi}(P||Q)=\sup_{g\in\mathcal{G}}V_{2}(g).$
where we have expressed the quantities to be maximized in Lemma 2.1 and Lemma 2.2 respectively as $V_{1}(g)$ and $V_{2}(g)$. It is clear that for each specific $g\in\mathcal{G}$, we have $V_{2}(g)\geq V_{1}(g)$ (since Lemma 2.1 is a special case of Lemma 2.2 where $\alpha=0$). This is when we say $V_{2}$ is a (point-wise) tighter variational representation than $V_{1}$.
In the two optimization problems above, the optimal solution in both cases gives rise to same divergence. That is, when the two optimization problems can be solved perfectly, there is no advantage of one representation over the other. However, when the global optimal of the optimization problems is not attainable, any $g$ obtained in an optimization effort gives rise to an estimate of the divergence, namely as $V_{1}(g)$ in the first case and as $V_{2}(g)$ in the second. It is then clear that due to $V_{2}$ being a tighter variational representation, $V_{2}$ consistently provides a closer approximation to $\mathrm{D}_{\phi}(P||Q)$ than $V_{1}$ for the same $g$.
This virtue carries over to the min-max training strategy of UDA. In that case, $g$ is reparameterized by $t\ell\circ h^{\prime}$. We desire the inner maximization, whether of $V_{2}$ or of $V_{1}$, to give rise to a good estimate of a similar divergence while the outer minimization aims to reduce this divergence. If the inner maximization is perfect, then the choice between $V_{1}$ and $V_{2}$ makes no difference. However, if the inner maximization isn’t perfect, $V_{2}$ gives a point-wise superior result compared to $V_{1}$, leading to a flatter optimization region around the optimal solution. In other words, using $V_{2}$ as the maximization objective gives a better estimate of the corresponding $f$-divergence than $V_{1}$.
@@ Appendix D Omitted Proofs and Additional Results
where Eq. (11) and (13) are by the triangle property of loss function, Eq. (12) is by Lemma B.2, and the last inequality is by the definition of $f$-DD. This completes the proof. ∎
@@
$R_{\nu}(h)\leq$ $R_{\mu}(h)+\widetilde{\mathrm{D}}^{h,\mathcal{H}}_{\phi}(\mu||\nu)+\lambda^{*}$
$\leq$ $R_{\hat{\mu}}(h)+2\hat{\mathfrak{R}}_{\mathcal{S}}(\mathcal{H}^{\ell})+\mathcal{O}\left(\sqrt{\frac{\log(1/\delta)}{n}}\right)+\widetilde{\mathrm{D}}^{h,\mathcal{H}}_{\phi}(\mu||\nu)+{\lambda}^{*}$ (18)
@@
where Eq. (18) and Eq. (19) are by Lemma B.3 and Lemma D.1, respectively. For the last inequality, notice that the loss is bounded between $[0,1]$ so the exponential function is $e$-Lipschitz in this domain, then we apply the Talagrand’s lemma [70, Lemma 5.7], i.e. ^$\hat{\mathfrak{R}}_{\mathcal{T}}(\exp\circ\mathcal{H}^{\ell})\leq e\hat{\mathfrak{R}}_{\mathcal{T}}(\mathcal{H}^{\ell})$.
Finally, due to the boundedness of the loss, there always exists a constant $\beta\in[1,e]$ such that $\beta\leq\min\left\{{{\mathbb{E}_{\nu}}\left[e^{\ell(h,h^{\prime})}\right]},{{\mathbb{E}_{\hat{\nu}}}\left[e^{\ell(h,h^{\prime})}\right]}\right\}$, and a simple choice is $\beta=1$. Plugging $\beta=1$ into Eq. (20) will complete the proof. ∎
D.3 Joint Error-Free Target Error Bound
[33] presents a $\lambda^{*}$-free target error bound by using the cross-domain error $\min\{R_{\nu}(f_{\mu}),R_{\mu}(f_{\nu})\}$. We now illustrate that it is possible to incorporate this term into all of our target error bounds.
In particular, the $\lambda^{*}$ term appears due to the following triangle inequalities:
$=$ $\underbrace{{{\mathbb{E}_{\nu}}\left[\ell(h,h^{*})\right]}-{{\mathbb{E}_{\mu}}\left[\ell(h,h^{*})\right]}}_{A_{1}}+\lambda^{*}.$
To avoid $\lambda^{*}$, we use $f_{\mu}$ and $f_{\nu}$ instead as the middle hypothesis for the application of triangle inequalities:
$R_{\nu}(h)-R_{\mu}(h)\leq$ $\underbrace{{{\mathbb{E}_{\nu}}\left[\ell(h,f_{\nu})\right]}-{{\mathbb{E}_{\mu}}\left[\ell(h,f_{\nu})\right]}}_{A_{2}}+{{\mathbb{E}_{\mu}}\left[\ell(f_{\nu},f_{\mu})\right]},$ (21)
$R_{\nu}(h)-R_{\mu}(h)\leq$ $\underbrace{{{\mathbb{E}_{\nu}}\left[\ell(h,f_{\mu})\right]}-{{\mathbb{E}_{\mu}}\left[\ell(h,f_{\mu})\right]}}_{A_{3}}+{{\mathbb{E}_{\nu}}\left[\ell(f_{\mu},f_{\nu})\right]}.$ (22)
If $|A_{2}|$ and $|A_{3}|$ are upper bounded by some hypothesis class-specific divergence $d_{\mathcal{H}}(\mu,\nu)$, then combining Eq (21) and Eq. (22) will give us:
$R_{\nu}(h)\leq R_{\mu}(h)+d_{\mathcal{H}}(\mu,\nu)+\min\{R_{\nu}(f_{\mu}),R_{\mu}(f_{\nu})\},$
where $R_{\nu}(f_{\mu})={{\mathbb{E}_{\nu}}\left[\ell(f_{\mu},f_{\nu})\right]}$ and $R_{\mu}(f_{\nu})={{\mathbb{E}_{\mu}}\left[\ell(f_{\nu},f_{\mu})\right]}$. Clearly, if $\mathcal{H}$ is large enough such that $f_{\mu}$ and $f_{\nu}$ are all achievable, $|A_{1}|$, $|A_{2}|$ and $|A_{3}|$ are all upper bounded by our $f$-DD defined in this paper. In fact, as long as $f_{\mu}$ and $f_{\nu}$ are close enough to $\mathcal{H}$, our $f$-DD can upper bound $|A_{2}|$ and $|A_{3}|$, with similar adaptations used in [33, Lemma 4.1].
D.4 Proof of Lemma 4.1
@@ D.3 Joint Error-Free Target Error Bound
We hint that with more careful handling of the aforementioned step, one can obtain a bound for the absolute mean deviation (i.e. $\left|A\right|$) in the end. However, the current development suffices for our immediate objectives.
@@
$R_{\nu}(h)-R_{\mu}(h)\leq\inf_{t\geq 0}\frac{1}{t}\left(K_{\mu}(t)+\mathrm{D}^{h,\mathcal{H}}_{\phi}(\nu||\mu)\right)+\lambda^{*}.$
@@
$R_{\nu}(h)-R_{\mu}(h)\leq$ $\inf_{t}\frac{t}{2\phi^{\prime\prime}(1)}+\frac{\mathrm{D}^{h,\mathcal{H}}_{\phi}(\nu||\mu)}{t}+\lambda^{*}=\sqrt{\frac{2\mathrm{D}^{h,\mathcal{H}}_{\phi}(\nu||\mu)}{\phi^{\prime\prime}(1)}}+\lambda^{*}.$
@@
D.7 Proof of Lemma 4.2 and Generalization Bounds for $f$-DD
@@
$=$ $\sup_{h^{\prime}}\sup_{\alpha}{{{\mathbb{E}_{\nu}}\left[t_{0}\ell(h,h^{\prime})\right]}+{{\mathbb{E}_{\mu}}\left[\alpha-\phi^{*}(t_{0}\ell(h,h^{\prime})+\alpha)\right]}}-\left(\sup_{h^{\prime},t}\sup_{\alpha}{{\mathbb{E}_{\hat{\nu}}}\left[t\ell(h,h^{\prime})\right]}+{{\mathbb{E}_{\hat{\mu}}}\left[\alpha-\phi^{*}(t\ell(h,h^{\prime})+\alpha)\right]}\right)$
$\leq$ $\sup_{h^{\prime}}\sup_{\alpha}{{{\mathbb{E}_{\nu}}\left[t_{0}\ell(h,h^{\prime})\right]}+{{\mathbb{E}_{\mu}}\left[\alpha-\phi^{*}(t_{0}\ell(h,h^{\prime})+\alpha)\right]}-\left({{\mathbb{E}_{\hat{\nu}}}\left[t_{0}\ell(h,h^{\prime})\right]}+{{\mathbb{E}_{\hat{\mu}}}\left[\alpha-\phi^{*}(t_{0}\ell(h,h^{\prime})+\alpha)\right]}\right)}$
$\leq$ $\sup_{h^{\prime}}\left|{{\mathbb{E}_{\nu}}\left[t_{0}\ell(h,h^{\prime})\right]}-{{\mathbb{E}_{\hat{\nu}}}\left[t_{0}\ell(h,h^{\prime})\right]}\right|+\sup_{h^{\prime}}\sup_{\alpha}\left|{{\mathbb{E}_{\mu}}\left[\phi^{*}(t_{0}\ell(h,h^{\prime})+\alpha)\right]}-{{\mathbb{E}_{\hat{\mu}}}\left[\phi^{*}(t_{0}\ell(h,h^{\prime})+\alpha)\right]}\right|$
@@
where the last two inequalities are by the scaling property of Rademacher complexity and Talagrand’s lemma [70]. This completes the proof. ∎
@@
Under the conditions in Lemma 4.2. Let $\phi$ be twice differentiable and $\phi^{\prime\prime}$ is monotone, for any $h\in\mathcal{H}$, with probability at least $1-\delta$, we have
@@
Then, recall that $\ell\leq 1$, and we use Lemma B.5444Note that, in this context, the random variable has a non-zero mean. Therefore, the first expectation term in Eq. (6) within the proof of Lemma B.5 should be retained. to obtain that
@@
$=$ $B(t)t^{2}{{\mathbb{E}_{\mu}}\left[(g(X))^{2}\right]}+B(t)t^{2}(C^{\prime 2}-2C^{\prime})\left({{\mathbb{E}_{\mu}}\left[g(X)\right]}\right)^{2}-t(C^{\prime}-1){{\mathbb{E}_{\mu}}\left[g(X)\right]}$
$=$ $B(t)t^{2}{{\mathbb{E}_{\mu}}\left[(g(X))^{2}\right]}+B(t)t^{2}(C^{2}-1)\left({{\mathbb{E}_{\mu}}\left[g(X)\right]}\right)^{2}-tC{{\mathbb{E}_{\mu}}\left[g(X)\right]},$ (26)
@@
Since $0\leq{{\mathbb{E}_{\mu}}\left[g(X)\right]}\leq\min\{r_{1}+r,1\}$ and $g(\cdot)\in[0,1]$, we have
$B(t)t{{\mathbb{E}_{\mu}}\left[(g(X))^{2}\right]}+B(t)t(C^{2}-1)\left({{\mathbb{E}_{\mu}}\left[g(X)\right]}\right)^{2}-C{{\mathbb{E}_{\mu}}\left[g(X)\right]}$
$\leq$ $B(t)t{{\mathbb{E}_{\mu}}\left[g(X)\right]}+B(t)t(C^{2}-1)\min\{r_{1}+r,1\}{{\mathbb{E}_{\mu}}\left[g(X)\right]}-C{{\mathbb{E}_{\mu}}\left[g(X)\right]}.$
@@
$B(t)t{{\mathbb{E}_{\mu}}\left[g(X)\right]}+B(t)t(C^{2}-1)\min\{r_{1}+r,1\}{{\mathbb{E}_{\mu}}\left[g(X)\right]}-C{{\mathbb{E}_{\mu}}\left[g(X)\right]}\leq 0.$
@@
$\left(e^{t}-t-1\right)\left(1-\min\{r_{1}+r,1\}+C^{2}\min\{r_{1}+r,1\}\right)-tC\leq 0,$ (27)
@@
$R_{\nu}(h)\leq R_{\hat{\mu}}$ $(h)+\inf_{C_{1},C_{2}}\frac{\mathrm{D}^{h,\mathcal{H}_{r}}_{\rm KL}(\hat{\nu}||\hat{\mu})}{C_{1}}+C_{2}R^{r}_{\mu}(h)+\mathcal{O}\left(\hat{\mathfrak{R}}_{\mathcal{S}}\left(\mathcal{H}^{\ell}_{r_{1}}\right)+\hat{\mathfrak{R}}_{\mathcal{S}}\left(\mathcal{H}^{\ell}_{r}\right)+\hat{\mathfrak{R}}_{\mathcal{T}}\left(\mathcal{H}^{\ell}_{r}\right)\right)$
$+\mathcal{O}\left(\frac{\log(1/\delta)}{n}+\frac{\log(1/\delta)}{m}\right)+\mathcal{O}\left(\sqrt{\frac{r_{1}\log(1/\delta)}{n}}+\sqrt{\frac{r\log(1/\delta)}{n}}+\sqrt{\frac{r\log(1/\delta)}{m}}\right)+\lambda_{r}^{*}.$
@@
D.11 Generalization Bounds based on $\chi^{2}$-DD